OpenStudy (barreraa):

Evaluate -8(x- 3y) when x = -2 and y = -4 (Points : 1) 112 80 -80 -16

OpenStudy (anonymous):

This one is a little tricky with all the negatives. xD -8(x - 3y) We have to plug x and y in. -8(-2 - 3(-4)) Let's do what's in parenthesis first. (-2 - 3(-4)) -3 * -4 = 12 (2 negatives = a positive number) -2 + 12 = 10 Now, we have: -8(10) -8 * 10 = -80 <--- Answer. (: